# Monitoring of water volume in a porous reservoir using seismic data: Validation of a numerical model with a field experiment

### v3

@@ Acknowledgements
Appendix A Measured data, water table level -36.2 cm
In this section, the data is shown for all measurements from drop height 5 cm at the water table level -36.2 cm. Notably, the first measurement within this series led to biased water volume estimate, as elaborated in Section 5.1.
Figure 9: The figure depicts measured data traces, with the top panels displaying results in the time domain and the bottom panels showing the corresponding data converted into the frequency domain. These measurements were taken at a water table level of -36.2 cm. Specifically, the figure presents the initial measurement out of a series of three, all conducted from a drop height of 5 cm.

### v4

@@ Acknowledgements
Appendix A Measured data, water-table level -36.2 cm
In this section, the data is shown for all measurements from drop height 5 cm at the water-table level -36.2 cm. Notably, the first measurement within this series led to biased water volume estimate, as elaborated in Section 5.1.
Figure 10: The figure depicts measured data traces, with the top panels displaying results in the time domain and the bottom panels showing the corresponding data converted into the frequency domain. These measurements were taken at a water-table level of -36.2 cm. Specifically, the figure presents the initial measurement out of a series of three, all conducted from a drop height of 5 cm.